OpenStudy (anonymous):

Evaluate Log e (2+J5) giving the answer in the form a+jb .

OpenStudy (dumbcow):

1.68 + j1.19

OpenStudy (anonymous):

you looking for this?\[ \ln(2+J5)\]

OpenStudy (dumbcow):

yeah thats what i assumed

OpenStudy (anonymous):

so how did you get the answer..

OpenStudy (dumbcow):

haha calculator :)

OpenStudy (anonymous):

ln2+...

OpenStudy (anonymous):

how did you use calculater..

OpenStudy (anonymous):

you know that right.. \[\ln(a+b)\neq lna+lnb\]

OpenStudy (dumbcow):

well you would have to have a graphing calculator (TI 83 or better) ln(2+5i) , there is symbol for an imaginary number

OpenStudy (dumbcow):

yes i am aware of that

OpenStudy (anonymous):

but it is not i, it is j

OpenStudy (anonymous):

never mind then..

OpenStudy (dumbcow):

same thing, different notation for a complex number which has real part and imaginary part

OpenStudy (anonymous):

I've never seen a+jb http://en.wikipedia.org/wiki/Complex_number

OpenStudy (anonymous):

i assume this is a complex number

OpenStudy (anonymous):

usually written as \[\ln(a+bi)\] or more usually \[\log(a+bi)\]

OpenStudy (anonymous):

find it via \[\log(a+bi)=\ln(|a+bi|)+i\theta\] where \[\tan(\theta)=\frac{b}{a}\] and \[|a+bi|=\sqrt{a^2+b^2}\]

OpenStudy (anonymous):

I like this one

OpenStudy (anonymous):

can we say that then..\[|a+bi|=\sqrt{x^2+b^2}=\sqrt{29}\] \[\tan^{-1}( 5/2)=68=\theta\] \[\log(2+5i)=\ln|2+5i|+i \theta\] =\[\ln \sqrt{29}+i68=1.68+68i\]

OpenStudy (anonymous):

actually thisone is better\[\log(a+bi)=\log(|a+bi|)+i\theta\quad or\quad \ln(a+bi)=\ln(|a+bi|)+i\theta\]

OpenStudy (anonymous):

68/180=R/pi \[68 \quad degree \approx \frac25\pi\] \[\log_{e} (2+5i)=\ln(2+5i)=1.68+\frac25\pi+2k \pi i\]